### v3

# YOLO-based Bearing Fault Diagnosis With Continuous Wavelet Transform

Abstract
This letter presents a locality-aware bearing fault diagnosis framework that operates on time-frequency representations and enables spatially interpretable decision-making. One-dimensional vibration signals are first mapped to two-dimensional time-frequency spectrograms using the continuous wavelet transform (CWT) with Morlet wavelets to enhance transient fault signatures. The diagnosis task is then formulated as object detection on the time-frequency plane, where YOLOv9, YOLOv10, and YOLOv11 are employed to localize fault-relevant regions and classify fault types simultaneously. Experiments on three public benchmarks, including Case Western Reserve University (CWRU), Paderborn University (PU), and Intelligent Maintenance System (IMS), demonstrate strong cross-dataset generalization compared with a representative MCNN-LSTM baseline. In particular, YOLOv11 achieves mAP@0.5 of 99.0% (CWRU), 97.8% (PU), and 99.5% (IMS), while providing region-aware visualization of fault patterns in the time-frequency domain. These results suggest that detection-based inference on CWT spectrograms provides an effective and interpretable complementary approach to conventional global classification for rotating machinery condition monitoring.
Index Terms:
Bearing fault diagnosis, continuous wavelet transform (CWT), time-frequency spectrogram, YOLO object detection.
I Introduction
Rolling bearings are critical components in rotating machinery, and their operational reliability directly influences equipment lifespan, production efficiency, and safety [1]. Prior studies report that nearly 40% of rotating machinery failures originate from bearing faults [2]. These faults are often embedded in dynamic, non-stationary vibration signals, making their early-stage diagnosis particularly challenging. Vibration signals acquired from rotating machinery are typically recorded as one-dimensional time-domain sequences. These raw signals contain temporal patterns related to mechanical faults, such as impulsive features caused by inner-race or ball defects [3]. However, these characteristics are often masked by noise and operational variability, especially during early-stage degradation [1]. Moreover, time-domain signals lack explicit frequency information, making it difficult to distinguish fault types that manifest at similar time intervals but differ in spectral behavior [2] (e.g., ball vs. outer-race faults, inner-race faults vs. imbalance, and looseness vs. misalignment).
Several traditional signal processing methods, such as short-time Fourier transform (STFT), Wigner-Ville distribution (WVD) [4], and Hilbert-Huang transform (HHT) [5], have been widely used to characterize non-stationary signals in both time and frequency domains. However, each of these methods exhibits intrinsic limitations [1, 2]. STFT employs a fixed analysis window, resulting in a fundamental trade-off between time and frequency resolution, which restricts its ability to capture features across multiple scales. WVD offers high resolution in theory, but it suffers from severe cross-term interference when analyzing multi-component signals. This interference often leads to spurious artifacts in the resulting spectrograms, which complicates the interpretation of fault features. HHT is data-adaptive and can decompose signals into intrinsic mode functions (IMFs), but its performance is highly sensitive to noise. Furthermore, it lacks a rigorous mathematical foundation for its empirical mode decomposition (EMD), which undermines the stability and reliability of its instantaneous frequency representation.
In contrast, continuous wavelet transform (CWT) [6], particularly with Morlet wavelets, offers multi-resolution analysis with better localization of both time and frequency content. Its ability to highlight transient features and scale-varying patterns makes it especially effective for capturing early bearing faults [7], which may be subtle and masked by background variations. Therefore, CWT serves as a powerful front-end transformation for fault-related pattern enhancement in complex industrial settings.
While CWT effectively transforms vibration signals into informative two-dimensional time-frequency representations, the subsequent classification stage remains critical. In recent years, deep learning (DL) methods have emerged as powerful tools for bearing fault diagnosis due to their end-to-end learning capability and feature extraction strength [3, 8, 1, 2]. Xia et al. [3] demonstrated early success using CNNs on multi-sensor signals. Pan et al. [8] proposed a hybrid CNN-LSTM model, laying the foundation for Chen et al. [9], who further extended it with multi-scale kernels and stacked LSTMs, and reported over $98\%$ accuracy on bearing datasets.
Despite these achievements, most DL-based models perform global classification on entire signal segments, lacking the spatial interpretability and localization capabilities necessary for fault-aware maintenance. Hakim et al. [10] emphasized this limitation in their systematic taxonomy, highlighting challenges such as interpretability, data dependency, and the need for lightweight architectures suitable for embedded deployment.
To address these limitations, we propose a detection-based diagnosis framework that explicitly exploits locality in the time-frequency plane. Specifically, we first apply Morlet-based CWT to convert vibration segments into time-frequency spectrograms, where fault-related transients manifest as spatially localized energy patterns. We then cast bearing diagnosis as an object detection problem on these spectrograms and adopt YOLO-based detectors, including YOLOv9 [11], YOLOv10 [12], and YOLOv11 [13], to jointly localize fault-relevant regions and classify fault types. Compared with conventional global classification, the proposed formulation enables region-aware interpretability and can be more robust to operating variations because decisions are made based on localized time-frequency evidence. To the best of our knowledge, this is the first study to systematically investigate modern YOLO variants (v9–v11) for spectrogram-level fault localization and diagnosis across multiple public bearing benchmarks.
The main contributions of this letter are as follows:
  • We introduce a locality-aware diagnosis pipeline that combines Morlet-based CWT with detection-based inference on time-frequency representations for bearing fault analysis.
  • We reformulate bearing fault diagnosis from global segment classification to object detection on the time-frequency plane, enabling spatially interpretable localization of fault-relevant regions.
  • We provide an extensive evaluation on three public benchmarks, including CWRU [14], PU [15], and IMS [16], demonstrating strong generalization under diverse operating conditions and fault types.
  • We benchmark three modern YOLO variants (YOLOv9–YOLOv11) under a unified spectrogram-based detection setting, highlighting the accuracy and efficiency trade-off for deployment-oriented condition monitoring.
II System Overview and Proposed CWT-YOLO
The proposed system consists of three stages: (1) CWT-based time-frequency transformation, (2) spectrogram pre-processing and bounding-box annotation, and (3) YOLO-based detection for joint fault localization and classification. Fig. 1 shows representative vibration segments under different bearing conditions, where discriminative transients are often difficult to identify directly in the time domain.
Figure 1: Sample vibration signals for four bearing conditions: (a) Normal, (b) Ball fault, (c) Inner race fault, (d) Outer race fault. Time (x-axis) and amplitude (y-axis) are shown for illustration, motivating a time-frequency transformation to expose discriminative fault signatures.
II-A Continuous Wavelet Transform (CWT)
The CWT is a time-frequency analysis tool that decomposes a one-dimensional signal into localized time-scale components using a family of wavelets. Compared to fixed-window methods like the STFT, CWT enables multi-resolution analysis with dynamic windowing, making it especially suitable for analyzing non-stationary signals such as bearing vibrations. The CWT of a signal $x(t)$ is defined as:
$\text{CWT}(a,b)=\frac{1}{\sqrt{|a|}}\int_{-\infty}^{\infty}x(t)\,\psi^{*}\left(\frac{t-b}{a}\right)dt,$ (1)
where $a$ denotes the scale (inverse frequency), $b$ is the time shift, $\psi(t)$ is the mother wavelet, and $\psi^{*}(t)$ its complex conjugate.
In this work, we adopt the Morlet wavelet due to its favorable time-frequency localization and robustness to high-frequency noise. The Morlet wavelet combines a sinusoidal carrier with a Gaussian envelope, providing favorable time-frequency localization and robustness to high-frequency noise. In our implementation, we use a real-valued Morlet form $\psi(t)=e^{-\frac{t^{2}}{2}}\cos(5t)$, which offers balanced resolution across different frequency bands and has shown strong performance for early fault detection [7]. By applying CWT to raw vibration signals, we generate two-dimensional time-frequency spectrograms that expose transient features and scale-dependent energy distributions. These spectrograms serve as spatially structured inputs for subsequent object detection models.
II-B Time-Frequency Spectrogram Visualization
To validate the visual enhancement provided by CWT, we convert the same vibration signals from Fig. 1 into two-dimensional time-frequency spectrograms using Morlet-based wavelet transformation. The resulting images, shown in Fig. 2, capture distinct energy concentration patterns that correspond to different fault types.
Compared to the original one-dimensional signals, these spectrograms exhibit spatially localized frequency bursts that are visually separable. For example, inner race and ball faults typically produce transient, broadband responses, whereas outer race faults show localized energy around characteristic frequencies. These patterns emerge clearly across the time and scale axes after CWT, forming structured visual cues that can be exploited by object detection models. This pre-processing not only standardizes the visual data across datasets but also improves feature discrimination and training stability.
Figure 2: CWT-based spectrograms for four bearing conditions: (a) Normal, (b) Ball fault, (c) Inner race fault, (d) Outer race fault.
II-C Data Pre-processing and Labeling
To ensure that the CWT-generated spectrograms are suitable for training YOLO-based object detectors, a series of pre-processing and annotation steps are applied to the vibration data prior to model inference.
II-C1 Signal Segmentation and Spectrogram Generation
For each dataset (CWRU, PU, and IMS), the raw vibration signals are segmented into fixed-length windows, each containing a sufficient number of samples to capture fault-relevant transients. Empirically, a window length of $2048$ samples with $50\%$ overlap is used to balance time resolution and data diversity. Each segment is transformed into a 2D time-frequency spectrogram using the Morlet-based CWT described earlier. The resulting grayscale spectrograms are logarithmically compressed and normalized to a $[0,1]$ range.
All spectrograms are resized to $640\times 640$ pixels to meet the input specification of YOLOv9, v10, and v11 models. This image resolution provides a balance between spatial detail preservation and computational efficiency during training and inference. The resolution of $640\times 640$ was empirically selected to balance spectral resolution and GPU memory constraints, while aligning with YOLO model input requirements.
II-C2 Bounding Box Annotation for Object Detection
Unlike traditional classification models that produce a single label for an entire signal segment, YOLO models require localized object annotations in the form of bounding boxes. To enable this, we manually annotate each spectrogram using the LabelImg tool, enclosing the time-frequency regions that exhibit fault-specific energy concentrations. Each box is associated with a class label corresponding to one of four fault types: Normal, Ball fault, Inner race fault, and Outer race fault.
These annotations are saved in YOLO format, which specifies each object as a row of normalized values [class_id, x_center, y_center, width, height] with respect to the image dimensions. For consistency, each dataset is annotated separately to ensure accurate spatial alignment with the visual fault features. To improve reproducibility, we adopt a consistent annotation criterion across datasets. Bounding boxes are drawn to cover the most prominent time-frequency energy concentrations associated with each fault class, while excluding background regions with diffuse responses. When multiple transient blobs appear in a spectrogram, the box encloses the dominant cluster that is visually persistent across samples of the same class, which reduces subjective variation in ROI selection.
II-C3 Dataset Splitting and Augmentation
The labeled dataset is partitioned into training, validation, and test sets using an $80:10:10$ ratio to ensure sufficient coverage and generalization. To further enhance model robustness, we apply data augmentation techniques including horizontal flipping, small-scale rotation $\pm 5^{\circ}$, and contrast jittering. These augmentations simulate minor variations in vibration patterns and imaging conditions that may occur in real-world settings.
Through this pre-processing pipeline, the spectrogram data are transformed into a high-quality object detection dataset with spatially grounded annotations, enabling YOLO models to learn fault-localizing features in both time and frequency domains.
II-D YOLO-based Fault Detection
In this study, bearing fault diagnosis is formulated as an object detection problem on CWT-generated spectrograms. Rather than assigning a global label to each signal segment, the system identifies and localizes regions of interest (ROIs) where fault-related energy patterns occur, enabling both classification and spatial localization.
We adopt YOLOv9 [11], YOLOv10 [12], and YOLOv11 [13] for their strong detection performance and architectural enhancements tailored to lightweight applications. YOLOv9 introduces programmable gradient information (PGI) and the generalized efficient layer aggregation network (GELAN) backbone for improved gradient flow and multi-scale representation, while removing anchor boxes via an anchor-free detection head. YOLOv10 eliminates the need for non-maximum suppression (NMS) through a dual-label assignment strategy, and uses spatial-channel decoupled downsampling (SCDown) with rank-guided channel interaction block (CIB) blocks to reduce redundancy. YOLOv11 further streamlines the architecture with C3k2 (cross-stage convolutional module with kernel size 2) blocks, the C2PSA (cross-stage partial with spatial attention) attention mechanism, and maintains minimal floating-point operations (FLOPs) and parameter count, making it ideal for embedded deployment.
All models are trained on $640\times 640$ spectrograms derived from the CWRU [14], PU [15], and IMS [16] datasets, using bounding box annotations over energy-dense fault regions. A composite loss function including localization, objectness, and classification terms guides the training.
III Experimental Results
III-A Performance Metrics
We evaluate model performance using the following metrics:
  • Mean Average Precision (mAP): Computed by averaging the area under the precision-recall curve across all classes and all intersection-over-union (IoU) thresholds. In this study, mAP is reported at an IoU threshold of $0.5$ (denoted as mAP@0.5), following standard object detection evaluation protocols.
  • Precision (PRE): The proportion of correctly identified positive predictions is calculated by
    $\text{PRE}=\frac{TP}{TP+FP}.$ (2)
  • Recall (REC): The ability of the model to identify all relevant instances is calculated by
    $\text{REC}=\frac{TP}{TP+FN}.$ (3)
  • F1 Score: The harmonic mean of precision and recall is calculated by
    $\text{F1}=2\cdot\frac{\text{PRE}\cdot\text{REC}}{\text{PRE}+\text{REC}}.$ (4)
Here, TP (true positives) are correctly detected fault regions; FP (false positives) are normal regions misclassified as faults; FN (false negatives) are missed faults.
III-B Experimental Setup
All models are trained and evaluated on three public bearing fault datasets: CWRU [14], PU [15], and IMS [16]. The CWT spectrograms are generated from raw vibration signals with 2048-sample windows and $50\%$ overlap. Images are resized to $640\times 640$ pixels and annotated with bounding boxes using the LabelImg tool. The datasets are partitioned into 80% training, $10\%$ validation, and $10\%$ testing splits. All YOLO models are trained for $500$ epochs using stochastic gradient descent (SGD) (learning rate=0.01, batch size=8) on a workstation with an Intel Core i7-class CPU and an NVIDIA RTX 3070 Ti GPU. Unless otherwise stated, the same training protocol and data split are used for all YOLO variants to ensure a fair comparison.
III-C Performance Comparison Across Datasets
Table I reports the mean average precision (mAP@0.5), precision (PRE), recall (REC), and F1 score (F1) for each model on the three datasets. Overall, YOLO-based detectors consistently outperform the MCNN-LSTM baseline [9] across datasets, indicating that detection-based inference on time-frequency representations can better capture localized fault signatures. Notably, YOLOv11 achieves the strongest results on PU and competitive performance on CWRU and IMS, suggesting a favorable accuracy–efficiency trade-off among the evaluated variants. The strong PU performance of YOLOv11 can be partly explained by its attention mechanism (C2PSA), which helps emphasize localized patterns under varying operating conditions.
To compare computational complexity, Table II lists the estimated FLOPs and parameter counts (Params) under the adopted implementations and the same input resolution. Values for YOLO variants are computed using PyTorch on the hardware described in Section III-B. For MCNN-LSTM, the values correspond to its original compact 1D time-series architecture.
TABLE I: Performance (mAP@0.5, Precision, Recall, F1) on CWRU, PU, and IMS.
Dataset Model mAP@0.5 PRE REC F1
CWRU [14] YOLOv9 99.4% 98.6% 98.5% 98.6%
YOLOv10 99.4% 99.2% 98.1% 98.6%
YOLOv11 99.0% 93.9% 98.5% 96.2%
MCNN-LSTM [9] 96.0% 96.1% 96.1% 96.1%
PU [15] YOLOv9 91.6% 80.8% 84.8% 82.7%
YOLOv10 97.2% 89.0% 92.7% 90.8%
YOLOv11 97.8% 94.9% 93.8% 94.3%
MCNN-LSTM [9] 77.7% 77.7% 77.4% 77.6%
IMS [16] YOLOv9 99.5% 99.9% 100.0% 100.0%
YOLOv10 99.5% 99.9% 100.0% 99.9%
YOLOv11 99.5% 100.0% 100.0% 100.0%
MCNN-LSTM [9] 96.8% 96.8% 96.8% 96.8%
TABLE II: Comparison of model complexity (independent of dataset).
Model FLOPs (G) Params (M)
YOLOv9 236.7 48.35
YOLOv10 8.2 2.57
YOLOv11 6.3 2.46
MCNN-LSTM [9] 0.010 0.352
Compared with YOLOv9, YOLOv10 and YOLOv11 substantially reduce computational complexity, while maintaining competitive detection performance across datasets. Although YOLOv9 achieves slightly higher mAP on CWRU, YOLOv11 provides a favorable balance between accuracy and efficiency, and it performs particularly well on PU and IMS. These results indicate that detection-based inference on CWT spectrograms can generalize across datasets with diverse operating conditions, while enabling interpretable fault localization through region-aware predictions.
It is also worth noting that the complexity values in Table II are reported for reference under the adopted implementations and input resolution, and the complexity between 2D spectrogram-based detectors and 1D time-series classifiers is not directly comparable. Nevertheless, the proposed framework suggests a practical accuracy and interpretability gain when leveraging localized time-frequency evidence for bearing fault diagnosis.
IV Conclusion
This letter presents a locality-aware bearing fault diagnosis framework that combines Morlet-based CWT with YOLO-based object detection on time-frequency spectrograms. By casting diagnosis as detection on the time-frequency plane, the proposed approach enables localization-driven inference and provides region-level interpretability by highlighting fault-relevant patterns. Experiments on the CWRU, PU, and IMS datasets demonstrate that YOLO-based detectors substantially improve generalization compared with the MCNN-LSTM baseline, while YOLOv11 offers a favorable balance between detection performance and computational complexity. These findings suggest that spectrogram-level detection is a practical and interpretable complementary approach to conventional global classification for rotating machinery condition monitoring.